A car slows down from 25 m/s to rest in a distance of 93 m. What was its acceleration, assumed constant?
1
Expert's answer
2012-09-21T11:36:09-0400
Let's find out the deceleration time of a car:

V = at ==> t = V/a

We know that

L = at²/2,

so

a = 2L/t² = 2L/(V/a)² = 2La²/V²,

or

1 = 2La/V² ==> a = V²/(2L).

At last,

a = 25²/(2*93) ≈ 3.3602 m/s².

So, the acceleration of a car was approximately -3.3602 m/s².